I am in a course and one of the exaples is y=-2x+1.
I am having problems figuring out how to find the other 2 componets to make this please help. in a straight line.

Found 2 solutions by rfer, mananth:
Answer by rfer(16322)

make an input/output table
y=-2x+1
y=-2*0+1
y=1
(0,1)
----------
0=-2x+1
2x=1
x=1/2
(1/2,0)
---------------
y=-2*2+1
y=-4+1
y=-3
(2,-3)

Answer by mananth(16949)

y=-2x+1
Find the y & x intercept points
Y intercept plug x=0
when x= 0 y= 1 ( 0 , 1 )
x intercept point plug y=0
when y = 0 x= 0.5 ( 0.5 , 0 )
You have the two points . mark the points and draw a line joining the points.